F(1)=7, f(n)=-4xf(n-1)+15
Answer by stanbon(75887)

F(1)=7, f(n)=-4xf(n-1)+15
---------
F(2) = -4x7+15 = -13
---------
F(3) = -4x(-13)+15 = 52+15 = 67
---------
F(4) = -4x(67)+15 = -253
---------
etc.
-----
Cheers,
Stan H.
-----------